# thanks: Corresponding author

### v1

@@ III Projected ensemble and emergence of deep thermalization with $\tilde{M}$
Figure 2: (a)-(b) Evolution of $\overline{\Delta^{(1)}}$ with time for $\tilde{M}=0.2$ and $0.4$ respectively. (c)-(d) Evolution of $\overline{\Delta^{(2)}}$ with time for $\tilde{M}=0.2$ and $0.4$ respectively. The overbar suggests quantity averaged over graph realizations. All data represented here are averaged over $500$ graphs. Power-law fits for $L=14$ are shown with black dashed lines, and the corresponding exponents are noted in the plot. Vertical black dotted lines indicate the region where the fitting was done to obtain the power-law exponents. $L_{A}=2$. The parameters for Hamiltonian (Eq. 1) are taken as $g=1.0,h=0.1$.
First, we analyze the convergence of PE to Haar ensemble with time. Figure 2 shows the evolution of trace distance from Haar ensemble with time for the first $2$ moments of PE for multiple $\tilde{M}$. We take $L_{A}=2$ for the subsequent analysis in this section. $\overline{\Delta^{(1)}},\overline{\Delta^{(2)}}$ decays with time for all cases, ultimately saturating to a system size dependent asymptotic value $\overline{\Delta^{(k)}}(t\to\infty)$. The decay with time shows a power-law scaling with the exponent dependent on $\tilde{M}$. The convergence to Haar ensemble is faster in the chaotic regime ($\tilde{M}=0.4$), where it decays with time as $\propto t^{-1.3}$. The exponent decreases as $\tilde{M}\to 0,1$. The power-law exponent appears same for all the moments (upto the error in their estimation). This indicates that the system thermalizes at the same rate for all the moments.
@@
Figure 3: (a)-(c) Asymptotic($t=10^{9}J$) trace distance of first 3 moments of PE from Haar ensemble as a function of $L$. The asymptotic trace distance gets suppressed with increasing $L$ for all $\tilde{M}$. Error bars represent the standard error of mean. (d)-(f) The same data plotted to show the the scaling of $\overline{\Delta^{(k)}}(t\to\infty)$ with $L$ across all $\tilde{M}$ for first $3$ moments of PE. Each data point shown is an average over 500 graph realizations, calculated for a subsystem of size $L_{A}=2$ with Hamiltonian (Eq. 1) parameters $g=1.0$ and $h=0.1$.
Next, we focus on the trend of asymptotic(late-time) trace distance $\overline{\Delta^{(k)}}(t\to\infty)$ with $\tilde{M}$. We compare the asymptotic trace distance between the first three moments of PE and the corresponding moments of the Haar ensemble in Fig. 3(a)-(c) as a function of $\tilde{M}$ for multiple $L$. The asymptotic trace distance is large at extreme values of $\tilde{M}$ (where WD spectral statistics breaks down) and goes down at intermediate $\tilde{M}$ where the spectrum is expected to be chaotic. The results are obtained after averaging over $500$ graph realizations for each $\tilde{M}$. The trace distance has very weak dependence on $L$ at $\tilde{M}\to 0,1$ i.e., where the system approaches an integrable or localized limit. The $\Delta^{(k)}(t\to\infty)$ shows slight asymmetry with $\tilde{M}$ with a minimum appearing around $\tilde{M}\approx 0.5$. Finally, in Fig. 3(d)-(f), we present the scaling of $\overline{\Delta^{(k)}}(t\to\infty)$ with $L$. The asymptotic trace distance from the Haar ensemble decays exponentially with $L$ for intermediate $\tilde{M}$. This supports the convergence of PE to the Haar ensemble at the thermodynamic limit, consistent with the deep thermalization in generic chaotic models at $L\to\infty$ [38, 39]. The decay exponent with $L$ depends on $\tilde{M}$. The trace distance decays weakly with $L$ for $\tilde{M}\to 0,1$, while it shows the fastest decay with $L$ at intermediate $\tilde{M}$. At the fully connected limit, the trace distance is almost independent of system size. The qualitative behavior is the same for all three moments. The breakdown of deep thermalization $\tilde{M}\to 0,1$, can be attributed to the approximate conserved charges close to the integrable limit, which modifies the asymptotic (in $L,t$) ensemble away from the Haar ensemble. This reflects on $\Delta^{(1)}$ itself as the density matrix gets modified to a GGE from the Identity matrix. Note that the minimum of trace distance for a given $L,L_{A}$ appears slightly left of $M=0.5$. This slight asymmetry in asymptotic trace distance as a function of $\tilde{M}$ can be an artifact of small system sizes studied. Similar asymmetry has been observed in level spacing ratio in Ref. [32, 24]).
We have presented the data at the parity-broken ($h=0.1$) limit in this section. When parity symmetry is preserved, measurement outcomes should be post-selected to remain within a given parity sector [55, 51].
IV Dependence of partial SFF on $\tilde{M}$
@@
Having characterized the emergence of chaos through the properties of the state ensemble, we now turn to a complementary diagnostic: the correlations within the energy spectrum itself. To this end, we investigate the partial spectral form factor (pSFF), which generalizes the SFF to subsystems and serves as a probe of spectral correlation, a key signature of quantum chaos. Unlike the full SFF, the pSFF is more readily accessible in experiments. In this section, we utilize the pSFF to study the localization-to-chaos transition as a function of connectivity ($\tilde{M}$) and explore its dependence on both system and subsystem size.
@@
Before discussing the pSFF, it is instructive to review the SFF [60, 61, 18, 62] first. The SFF is a spectral diagnostic of quantum chaos as it captures both short and long-range spectral correlations, providing information beyond that contained in local level-spacing statistics (which captures only information of adjacent level spacings). Formally, SFF is the Fourier transform of the eigenvalue density-density correlation function, defined as follows,
$K(t)=\frac{1}{d^{2}}\sum_{i,j=1}^{d}e^{i(E_{i}-E_{j})t}=\frac{1}{d^{2}}\mathrm{Tr}[T(t)]\mathrm{Tr}[T^{\dagger}(t)]\ ,$ (6)
@@
where $d$ is the Hilbert space dimension of the full system and $E_{i}$ is $i$-th eigenvalue of the Hamiltonian. $t$ is a Fourier variable which can be taken as time. The time evolution unitary operator is represented as, $T(t)=e^{-\imath Ht}$. The prefactor ensures $K(0)=1$. In a chaotic spectra (obeying WD statistics), SFF decays from $1$ at $t=0$ till Thoules time $t_{Th}\approx O(\delta E^{-1})$ where $\delta E$ is the mean level spacing. This results in SFF dropping to below $1/d$ at intermediate time. This is followed by a linear ramp up to saturation to $1/d$ at plateau time (the time at which the plateau in SFF begins), $t_{p}\approx O(2^{L})$. The dip in SFF (referred to as a correlation hole) is caused by level repulsion and spectral rigidity in chaotic systems as predicted by RMT. This ramp-and-plateau structure is a signature of the chaotic spectra. For a spectrum following Poisson statistics, the SFF is expected to gradually reach its asymptotic value after an initial transient period without any correlation hole or ramp. For chaotic systems, the asymptotic SFF is given as $K(t\to\infty)=1/d$ in the absence of any degeneracy in the eigenspectrum. Recently, SFF has been experimentally measured in superconducting qubits using a randomized measurement protocol [40, 41, 63]. Ref [32] analyzes SFF in this model for $h=0$. In the intermediate regime of both $\tilde{M}$ and $g$, the SFF converges with increasing $L$ to the characteristic dip-ramp-plateau structure predicted by RMT, indicating chaotic spectra. Conversely, the chaos is suppressed at the extreme limits of connectivity. In both the sparsely connected ($\tilde{M}\to 0$) and densely connected ($\tilde{M}\to 1$) regimes, the development of the correlation hole is slower with increasing $L$. Similarly, the correlation hole vanishes for both very small and very large $g$, indicating a breakdown of RMT statistics.
@@
However, despite recent experimental demonstrations of SFF measurement on superconducting qubit arrays, a significant practical challenge remains regarding the requirement of global control and site-resolved operations across the entire quantum system. In practice, implementing such fully controlled rotations on all qubits becomes technically demanding and resource-intensive, particularly as system size increases. The partial spectral form factor (pSFF) was recently introduced to address this challenge [40], and unlike the full SFF, it requires controlled operations that scale only with the subsystem size [40, 41]. The pSFF is described in terms of the time-evolution operator on a subsystem. The pSFF can be defined as follows,
@@
$=\frac{1}{dd_{A}}\mathrm{Tr}_{B}[\mathrm{Tr}_{A}[T(t)]\,\mathrm{Tr}_{A}[T^{{\dagger}}(t)]],$ (7)
@@
where $\rho_{B}(E_{i})=\mathrm{Tr}_{A}\left[|E_{i}\rangle\langle E_{i}|\right]$ is the reduced density matrix of subsystem $B$ for energy eigenstate $|E_{i}\rangle$. pSFF captures correlation of both eigenstates and eigenvalues. Similar to SFF, the normalization factor in Eq.  IV ensures $K_{A}(t=0)=1$.
@@
Figure 4: Time evolution of the pSFF, $\overline{K_{A}(t)}$, for various system sizes ($L$, different colors) and subsystem sizes ($L_{A}$, different line styles). Panels show different connectivities: (a) $\tilde{M}=0.2$, (b) $\tilde{M}=0.4$, (c) $\tilde{M}=0.9$, and (d) $\tilde{M}=1.0$. The solid, dashed, and dotted lines correspond to $L_{A}=1,2,$ and $3$, respectively. The horizontal black lines denote the theoretical asymptotic values for a chaotic system, $1/d_{A}^{2}$. All data are averaged over $500$ graph realizations (denoted by overbar) with Hamiltonian (Eq. 1) parameters $g=1.0$ and $h=0.1$.
We now utilize the pSFF to analyze the Hamiltonian in Eq. 1, setting $g=1$ and $h=0.1$. For the following results, all quantities are averaged over $500$ graph realizations, and the pSFF is calculated on the entire spectrum without an unfolding procedure. We find that, similar to the full SFF, the pSFF universally exhibits an initial polynomial decay. Note that we exclude data from the sparse graphs ($\tilde{M}<0.3$), as the possibility of the subsystem sites belonging to disconnected clusters can cause persistent oscillations in pSFF. We analyze the behavior of pSFF at multiple values of connectivity, $\tilde{M}$, with the results presented in Fig. 4.
First, we describe the intermediate $\tilde{M}$ regime. pSFF shows a dip and a subsequent ramp-and-plateau structure. This correlation hole appears due to correlation in energy spectra, as expected in chaotic systems. The depth of the correlation hole depends on $L,L_{A}$. The depth increases with increasing $L$ and $L_{A}$. The time at the appearance of the correlation hole, $t_{Th}$ increases with $L$ for fixed $L_{A}$. For chaotic systems, RMT predicts [40] (in the absence of degeneracy) $K_{A}(t\to\infty)\approx K(t)+1/d_{A}^{2}=1/d+1/d_{A}^{2}$. For $L_{A}\ll L$, $K_{A}(t\to\infty)\approx 1/d_{A}^{2}$ i,e. asymptotic pSFF shows a subsystem size ($L_{A}$) dependent shift. In Fig. 4, these asymptotic values (for chaotic spectra) are shown using black horizontal lines. For $\tilde{M}=0.4$, Fig. 4(b) shows that the asymptotic $K_{A}$ shows very small fluctuation at late time and approaches $\approx 1/d_{A}^{2}$. Additionally, there is also no significant fluctuation at $t>t_{p}$ in pSFF when no average over graph realization is taken.
@@
In the sparsely connected regime, the pSFF shows qualitatively different behavior from the chaotic case. For $\tilde{M}=0.2$, Fig. 4(a) shows that there is no distinct correlation hole, and the pSFF reaches its asymptotic value after a transient period. The absence of a correlation hole indicates the lack of correlation in the eigenspectrum (consistent with Poisson statistics). There is also a persistent oscillation present at late times after averaging over graph realizations. This can be attributed to quasi-degeneracies in the eigenspectrum (when adjacent level spacing goes below $\delta E$) [32]. For $\tilde{M}$ close to 0, there is a significantly large fluctuation at late time for a single graph realization.
@@
In the densely connected limit approaching all-to-all connectivity, i.e., $\tilde{M}=0.9$, Fig. 4(c) shows a correlation hole in pSFF. In this case, the asymptotic pSFF shows small oscillation around a value larger than $1/d_{A}^{2}$. This deviation from $1/d_{A}^{2}$ is caused by approximate degeneracy in the energy spectrum close to $\tilde{M}=1$. Fig. 4(d) shows the pSFF at $\tilde{M}=1$. A persistent oscillation around at $t>t_{p}$ can be observed. The saturation value at $t>t_{p}$ shows a large deviation from $1/d_{A}^{2}$, hinting at the degeneracy in the spectrum.
@@
At $\tilde{M}=1.0$, i.e., where the model is expected to be almost integrable (in the presence of a small nonzero $h$), we observe a dip in pSFF at an intermediate time. The pSFF eventually saturates to a value larger than the RMT prediction for chaotic systems, suggesting deviation from chaotic spectra.
@@
Figure 5: (Quantitative analysis of the pSFF. (a) Scaling of the plateau time, $\log_{10}t_{p}$, as a function of system size $L$ for several connectance, $\tilde{M}$, with a fixed subsystem size of $L_{A}=2$. (b) The late-time subsystem purity, $\overline{p_{B}}$, as a function of $\tilde{M}$ for a fixed system size $L=12$ and varying subsystem sizes, $L_{A}$. All data are averaged over $500$ graph realizations with Hamiltonian (Eq. 1) parameters $g=1.0$ and $h=0.1$. Here overbar denotes annealed average.
@@
To further quantify the ramp-and-plateau behavior and its dependence on $L$ in the chaotic regime, we study the scaling of the plateau time, $t_{p}$ with $L$ for $L_{A}=2$. This time is determined numerically as the intersection of two linear fits: one to the logarithmic plot of the ramp ($t_{Th}<t<t_{p}$) and another to the late-time plateau ($t>t_{p}$). We observe a clear exponential scaling of $t_{p}$ with system size $L$, as shown in Fig. 5(a), with a slope that does not strongly depend on $\tilde{M}$ within the chaotic regime. The exponential scaling of the plateau time is a feature [40, 41] of the SFF that is inherited by the pSFF. Our data in Fig. 5(a) quantitatively supports this behavior.
@@
Figure 6: Normalized distribution of the asymptotic (late-time) pSFF, $\overline{K_{A}}(t\to\infty)$, computed over all possible two-site subsystems for different system sizes, $L$. Panel (a) shows the distribution in the chaotic regime, $\tilde{M}=0.6$. Panel (b) shows the distribution in the near-integrable regime, $\tilde{M}=0.9$. For all calculations, the subsystem size is fixed at $L_{A}=2$ and the Hamiltonian (Eq. 1) parameters are $g=1.0$ and $h=0.1$.
@@
Another related quantity to pSFF is subsystem purity, defined as $p_{B}=\mathrm{Tr}(\rho_{B}^{2})$ where $\rho_{B}$ is the reduced density matrix of subsystem $B$. This quantity captures the entanglement of subsystem with its complement and is related to subsystem Renyi entropy as $R_{2}=-\log(p_{B})$. The purity can be obtained from pSFF at late time as $p_{B}=\lim_{t_{p}/t\to 0}K_{A}(t)d_{A}$. Figure 5(b) shows $p_{B}$ as a function of $\tilde{M}$ for multiple $L_{A}$ and $L=12$. The purity is minimum at the intermediate $\tilde{M}$ i.e. subsystem $B$ is more mixed (higher entropy) in the chaotic regime compared to the integrable or localized regime. At both $\tilde{M}\to 0,1$, the purity goes up. The purity goes down with increasing $L_{A}$. The higher purity at $\tilde{M}\to 0,1$ is consistent with the expectation of lower entanglement in the localized or integrable phase compared to the chaotic phase. The purity of $B$ decreases (i.e. increase in entanglement of subsystems) with increasing $L_{A}$.

### v2

@@ III Projected ensemble and emergence of deep thermalization
Figure 5: (a)-(b) Evolution of $\overline{\Delta^{(1)}}$ with time for $\tilde{M}=0.2$ and $0.4$, respectively. (c)-(d) Evolution of $\overline{\Delta^{(2)}}$ with time for $\tilde{M}=0.2$ and $0.4$, respectively. All data represented here are averaged over $500$ graphs. Power-law fits for $L=14$ are shown with black dashed lines, and the corresponding exponents are noted in the plot. Vertical black dotted lines indicate the region where the fitting was done to obtain the power-law exponents. (e)-(f) The slope of trace distance with time across $\tilde{M}$ (with estimated error in slope indicated). $L_{A}=2$. The parameters for Hamiltonian (Eq. 1) are taken as $g=1.0,h=0.1$.
@@
Figure 6: (a)-(c) Asymptotic ($t=10^{9}J^{-1}$) trace distance of first 3 moments of PE from Haar ensemble as a function of $L_{B}$. The asymptotic trace distance gets suppressed with increasing $L_{B}$ for all $\tilde{M}$. Error bars represent the standard error of the mean across realizations at a fixed $t$. (d)-(f) The same data plotted to show the the scaling of $\overline{\Delta^{(k)}}(t\to\infty)$ with $L$ across all $\tilde{M}$ for first $3$ moments of PE. The error bars are not shown when they are comparable to the marker size. Each data point shown represents an average over 500 graph realizations, calculated for a subsystem of size $L_{A}=2$ with Hamiltonian parameters (Eq. 1) $g=1.0$ and $h=0.1$. The black dashed line shows the trace distance of an empirical ensemble of $500$ random states (sampled from Haar random states of system size $L$) from the corresponding moments of the Haar ensemble.
Convergence of PE to the Haar ensemble with time:
For intermediate $\tilde{M}$, the trace distance between the PE and the Haar ensemble decays as a power law in time, with an $\tilde{M}-$dependent exponent $\alpha$ (Fig. 5(a)–(d)). The convergence is fastest in the chaotic regime, with the largest exponent ($\alpha\approx 1.6$) occurring at $\tilde{M}\approx 0.3$ (Fig. 5(e)–(f)). For comparison, the nearest-neighbor mixed-field Ising model at a specific chaotic parameter point (where it is expected to be chaotic) exhibits an exponent of $\approx 1.2$ [31]. The exponent decreases on either side of $\tilde{M}\approx 0.3$, toward both the sparse and dense limits. The power-law exponent is consistent within error bars across both the first and second moments of the PE, indicating that the thermalization rate is independent of the moment order, a feature previously reported in Refs. [31, 66]. At extremal values ($\tilde{M}\leq 0.1$, $\tilde{M}\geq 0.8$), the trace distance exhibits a slower approach to its asymptotic value, accompanied by pronounced plateaus, suggesting that thermalization proceeds via distinct dynamical timescales in these regimes.
Asymptotic (late-time) trace distance, $\overline{\Delta^{(k)}}(t\to\infty)$:
We compare the late-time asymptotic trace distances $\Delta^{(k)}$ for $k=1,2,3$ in Fig. 6(a)-(c) as a function of $\tilde{M}$ for multiple $L$. Figure 6(d)-(f) show the $L$ dependence of $\overline{\Delta^{(k)}}(t\to\infty)$.
At fixed $L$, the asymptotic trace distance is largest at extremal $\tilde{M}$ and decreases toward intermediate connectance, reaching a minimum at $\tilde{M}\approx 0.4$. At $\tilde{M}=0.4$$0.5$, the trace distance approaches the same obtained for Haar random states of the same system size (black dashed lines in Fig. 6(d)–(f)). The larger trace distance at $\tilde{M}$ away from $0.4$ suggests that these Hamiltonians possess approximate conservation laws that impose structure on the eigenstates, preventing full ergodicity (as discussed in Sec. II). We find no evidence of slow time-dependent drift in the asymptotic region (Fig. 5(a)–(d)), confirming that the trace distance has genuinely converged rather than remaining in a slow transient. On the sparse side ($\tilde{M}\ll 0.4$), disconnected clusters give rise to local conserved quantities, accounting for some of the observed deviation from chaos. On the dense side ($\tilde{M}\gtrsim 0.4$), the approximate $SU(2)$ symmetry inherited from the near-LMG structure provides the relevant approximate conservation laws that impede PE convergence.
The asymptotic trace distance $\overline{\Delta^{(k)}}(t\to\infty)$ decays exponentially with $L$ at intermediate $\tilde{M}$, consistent with convergence of the PE to the Haar ensemble in the thermodynamic limit and with deep thermalization in chaotic models. In contrast, at $\tilde{M}\to 0,1$, the decay with $L$ is significantly slower, while the fastest decay occurs at intermediate $\tilde{M}$. This qualitative behavior is consistent across the first, second, and third moments of the PE. The fact that the trace distance decays with $L$ even at $\tilde{M}=0.1$ and $0.9$, albeit slowly, is tentative evidence that the system may be chaotic throughout, except at the integrable limits $\tilde{M}=0,1$, a conclusion consistent with Ref. [42].
So far in this section, we have investigated the PE’s approach to the Haar ensemble for time evolutions starting from a specific zero-energy expectation-value initial state (Eq. 11). At $\tilde{M}=1.0$, the Hamiltonian has a permutation symmetry which is preserved by this state, as a result of which the instantaneous state has finite overlaps with only a small number of energy eigenstates of the $\tilde{M}=1$ Hamiltonian. However, the results shown in Fig. 6 stay valid even if we consider other 0 energy expectation value initial states that lack permutation symmetry. In Appendix C, we show analogous results for initial direct product states with limited or no permutation symmetry, demonstrating the robustness of the results presented in Fig. 6.
In summary, we find that finite-system studies of the PE show a clearly faster onset of chaos (through convergence of the PE to the Haar ensemble) with increasing system size at intermediate connectivity $\tilde{M}\sim 0.4-0.6$. Away from this, the approach of PE to the Haar ensemble is slower, likely due to approximate conserved charges or symmetries. In the latter cases, in accessible systems, the asymptotic PE approaches the Haar ensemble as system size increases, indicating that the number of approximately conserved charges and symmetries becomes subextensive. Measurement of the projected ensemble and its higher moments has been demonstrated in quantum simulators for 1D lattice systems with up to $25$ qubits [31]. Extending such protocols to random-graph connectivity presents additional experimental challenges, particularly in terms of qubit connectivity and mid-circuit measurement overhead. Nevertheless, our finite-size results suggest that the trace distance of the PE from the Haar ensemble provides a clear and resolvable diagnostic of the chaotic crossover, capturing the $\tilde{M}-$dependent convergence rate and the moment-independence of the thermalization exponent. These signatures are visible at system sizes accessible to near-term quantum devices, making this a tractable target for experimental investigation.
IV Partial Spectral Form Factor
@@
Having characterized the emergence of chaos through the properties of the projected ensemble, we now turn to a complementary diagnostic: the correlations within the energy spectrum itself. To this end, we investigate the partial spectral form factor (pSFF), which generalizes the spectral form factor (SFF) to subsystems and serves as a probe of spectral correlation, a key signature of quantum chaos. Unlike the SFF, the pSFF is more readily accessible in experiments. In this section, we utilize the pSFF to study the localization-to-chaos crossover as a function of connectivity ($\tilde{M}$) and explore its dependence on both system and subsystem size in different connectivity regimes.
@@
Before discussing the pSFF, we briefly review the SFF [59, 81, 94, 9]. The SFF is a spectral diagnostic of quantum chaos that encapsulates both short- and long-range spectral correlations, providing information beyond that contained in level-spacing statistics (which captures only adjacent level spacings). The SFF is the Fourier transform of the eigenvalue density-density correlation, defined as follows,
$K(t)=\frac{1}{d^{2}}\sum_{i,j=1}^{d}e^{i(E_{i}-E_{j})t}=\frac{1}{d^{2}}\mathrm{Tr}[\mathcal{T}(t)]\mathrm{Tr}[\mathcal{T}^{\dagger}(t)]\ ,$ (14)
@@
where $d$ is the Hilbert space dimension of the full system and $E_{i}$ is $i$-th eigenvalue of the Hamiltonian. $t$ is the Fourier variable of energy, with units of time. The time evolution unitary operator is represented as, $\mathcal{T}(t)=e^{-\imath Ht}$. The prefactor ensures $K(0)=1$. In a chaotic spectrum (obeying WD statistics), SFF decays from $1$ at $t=0$ till Thouless time $t_{Th}\approx O(\delta E^{-1})$ where $\delta E$ is the mean level spacing. This results in SFF dropping to below $1/d$ at intermediate time. This is followed by a linear ramp up to saturation to $1/d$ at plateau time (the time at which the plateau in SFF begins), $t_{p}\approx O(2^{L})$. The dip in SFF (referred to as a correlation hole) is caused by level repulsion and spectral rigidity in chaotic systems as predicted by random matrix theory (RMT). This ramp-and-plateau structure is a signature of the chaotic spectrum. For a spectrum following Poisson statistics, the SFF is expected to gradually reach its asymptotic value after an initial transient period without any correlation hole or ramp. For chaotic systems, the asymptotic SFF is given as $K(t\to\infty)=1/d$ in the absence of any degeneracy in the eigenspectrum. Recently, SFF has been experimentally measured in a 5-qubit superconducting device using a randomized measurement protocol [34]. Ref. [42] analyzed SFF in the model studied here for $h=0$. In the intermediate regime of both $\tilde{M}$ and $g$, the SFF converges with increasing $L$ to the characteristic dip-ramp-plateau structure predicted by RMT, indicating chaotic spectra. Conversely, chaos is suppressed at the extreme limits of connectivity. In both the sparsely connected ($\tilde{M}\to 0$) and densely connected ($\tilde{M}\to 1$) regimes, the development of the correlation hole is slower with increasing $L$. Similarly, the correlation hole vanishes for both very small and very large $g$, indicating a breakdown of RMT statistics.
@@
Despite recent experimental demonstrations of SFF measurement on superconducting qubit arrays, a significant practical challenge remains regarding the requirement of global control and site-resolved operations across the entire quantum system. In practice, implementing such fully controlled rotations on all qubits becomes resource-intensive as system size increases. The partial spectral form factor (pSFF) was recently introduced as a scalable alternative [52], which, unlike the full SFF, requires controlled operations that scale only within a smaller subsystem [52, 34]. The pSFF is described in terms of the time-evolution operator restricted to a subsystem. The pSFF can be defined as follows,
@@
$=\frac{1}{dd_{A}}\mathrm{Tr}_{B}[\mathrm{Tr}_{A}[\mathcal{T}(t)]\,\mathrm{Tr}_{A}[\mathcal{T}^{\dagger}(t)]],$ (15)
@@
where $\rho_{B}(E_{i})=\mathrm{Tr}_{A}\left[|E_{i}\rangle\langle E_{i}|\right]$ is the reduced density matrix of subsystem $B$ for energy eigenstate $|E_{i}\rangle$. pSFF captures the correlation of both eigenstates and eigenvalues. The normalization factor in Eq. IV ensures $K_{A}(0)=1$.
@@
Figure 7: Time evolution of the pSFF, $\overline{K_{A}}(t)$, for various system sizes ($L$, different colors) and subsystem sizes ($L_{A}$, different line styles). Panels show different connectivities: (a) $\tilde{M}=0.2$, (b) $\tilde{M}=0.4$, (c) $\tilde{M}=0.9$, and (d) $\tilde{M}=1.0$. The solid, dashed, and dotted lines correspond to $L_{A}=1,2,$ and $3$, respectively. The horizontal black lines denote the expected asymptotic values for a chaotic system, $1/d_{A}^{2}$. All data are averaged over $500$ graph realizations (denoted by an overbar) with Hamiltonian parameters (Eq. 1) of $g=1.0$ and $h=0.1$.
We now investigate pSFF for the Hamiltonian in Eq. 1, setting $g=1$ and $h=0.1$ with the goal of understanding the manifestation of localization, integrability and chaos as well as the finite-size effect in these cases. The SFF can be computed directly from the eigenvalue spectrum, allowing spectral unfolding on mid-spectrum eigenvalues to remove non-universal level-density variations. However, in an experimental setting, the full eigenspectrum is usually inaccessible, precluding spectral unfolding. In the following, we therefore compute the pSFF on the full spectrum without unfolding. In spite of this, we find that the pSFF reveals clear chaotic behavior at intermediate connectivity and distinct deviations at extremal connectivity, making the pSFF another route to investigate the model in scalable experimental systems.
We analyze the behavior of pSFF at multiple values of $\tilde{M}$ with the results presented in Fig. 7. In the results presented here, the pSFF is averaged over $500$ ER graph realizations.
pSFF at small $\tilde{M}$ :
@@
For $\tilde{M}=0.2$, Fig. 7(a) shows that there is no distinct correlation hole, and the pSFF reaches its asymptotic value after a transient period. The absence of a correlation hole indicates the lack of correlation in the eigenspectrum (consistent with Poisson statistics). There is also a persistent oscillation present at late-times, even after averaging over graph realizations. The black lines in Fig. 7(a) shows the asymptotic values of pSFF for a chaotic Hamiltonian. Apart from the late-time fluctuation in pSFF, the mean asymptotic value is also larger compared to the expectation for a chaotic Hamiltonian. This can be attributed to quasi-degeneracies in the eigenspectrum (when adjacent level spacing goes below $\delta E$) [42]. For $\tilde{M}$ close to 0, there is a wide fluctuation at all times due to the existence of disconnected clusters, which obscures the possible existence of any correlation hole.
pSFF at intermediate $\tilde{M}$ :
The pSFF shows qualitatively different behavior in the chaotic regime. In Fig. 7(b), the pSFF shows a dip (correlation hole) and a subsequent ramp-and-plateau structure. This correlation hole arises from energy-spectrum correlations, as expected in chaotic systems. The depth of the correlation hole increases with both $L$ and $L_{A}$. The time at the appearance of the correlation hole, $t_{\rm Th}$, increases with $L$ for fixed $L_{A}$. For chaotic systems, RMT predicts [52] (in the absence of degeneracy) $K_{A}(t\to\infty)\approx K(t\to\infty)+1/d_{A}^{2}=1/d+1/d_{A}^{2}$. For $L_{A}\ll L$, $K_{A}(t\to\infty)\approx 1/d_{A}^{2}$ i.e., asymptotic pSFF shows a subsystem size ($L_{A}$) dependent shift. In Fig. 7, these asymptotic values (for chaotic spectra) are shown using black horizontal lines. For $\tilde{M}=0.4$, Fig. 7(b) shows that the asymptotic $K_{A}$ shows very little fluctuation at late-time and approaches $\approx 1/d_{A}^{2}$. Additionally, there is no significant fluctuation in pSFF at $t>t_{p}$ when no average over graph realizations is taken.
pSFF at large $\tilde{M}$ :
@@
In the densely connected limit approaching all-to-all connectivity, i.e., $\tilde{M}=0.9$, Fig. 7(c) shows a dip in pSFF. In this case, the asymptotic pSFF shows small oscillation around a plateau value larger than $1/d_{A}^{2}$. This deviation from $1/d_{A}^{2}$ is caused by approximate degeneracy in the energy spectrum close to $\tilde{M}=1$.
pSFF at $\tilde{M}=1.0$ :
@@
At $\tilde{M}=1.0$, i.e., where the model is expected to be almost integrable (in the presence of a small nonzero $h$), we observe a dip in pSFF at an intermediate time, a feature that persists even in the $h=0$ limit. The pSFF eventually saturates to a value exceeding the RMT prediction for chaotic systems, suggesting a deviation from WD spectra.
@@
Figure 8: Quantitative analysis of the pSFF. (a) Scaling of the plateau time, $\log_{10}t_{p}$, as a function of system size $L$ for several connectance, $\tilde{M}$, with a fixed subsystem size of $L_{A}=2$. (b) The late-time subsystem purity, $\overline{p_{B}}$, as a function of $\tilde{M}$ for a fixed system size $L=12$ and varying subsystem sizes, $L_{A}$. All data are averaged over $500$ graph realizations (denoted by an overbar) with Hamiltonian parameters (Eq. 1) of $g=1.0$ and $h=0.1$. Here overbar denotes annealed average.
Scaling of the plateau time :
@@
To further quantify the ramp-and-plateau behavior and its dependence on $L$ in the chaotic regime, we study the scaling of the plateau time, $t_{p}$ with $L$ for $L_{A}=2$. This time is determined numerically as the intersection of two linear fits: one to the logarithmic plot of the ramp ($t_{\rm Th}<t<t_{p}$) and another to the late-time plateau ($t>t_{p}$). We observe a clear exponential scaling of $t_{p}$ with system size $L$, as shown in Fig. 8(a), with a slope that does not strongly depend on $\tilde{M}$ within the chaotic regime. The exponential scaling of the plateau time is a feature [52, 34] of the SFF that is inherited by the pSFF. Our data in Fig. 8(a) quantitatively supports this behavior.
@@
Figure 9: Normalized distribution of the asymptotic (late-time) pSFF, $K_{A}(t\to\infty)$, computed over all possible two-site subsystems for different system sizes, $L$. Panel (a) shows the distribution in the chaotic regime, $\tilde{M}=0.6$. Panel (b) shows the distribution in the near-integrable regime, $\tilde{M}=0.9$. For all calculations, the subsystem size is fixed at $L_{A}=2$ and the Hamiltonian (Eq. 1) parameters are $g=1.0$ and $h=0.1$. The data is presented here for $200$ graph realizations.
Entropy from pSFF :
@@
Another quantity related to pSFF is the purity of the subsystem averaged over eigenstates, defined as $p_{B}=\mathbb{E}_{i}[\mathrm{Tr}(\rho_{B}(i)^{2})]$ where $\rho_{B}(i)$ is the reduced density matrix of the subsystem $B$ in the $i$-th eigenstate. This quantity captures the entanglement of subsystem with its complement and is related to the annealed subsystem Renyi entropy as $R_{2}=-\log(p_{B})$. The purity can be obtained from pSFF at $t\to\infty$ as $p_{B}=\lim_{t_{p}/t\to 0}K_{A}(t)d_{A}$. Figure 8(b) shows $p_{B}$ as a function of $\tilde{M}$ for multiple $L_{A}$ and $L=12$. The purity is minimum at the intermediate $\tilde{M}$ i.e., subsystem $B$ is more mixed (more entanglement with $A$) in the chaotic regime compared to the integrable or localized regime. At both $\tilde{M}\to 0,1$, the purity goes up. The higher purity at $\tilde{M}\to 0,1$ is consistent with the expectation of lower entanglement in the localized or integrable eigenstates compared to the chaotic eigenstates. The purity of $B$ decreases (i.e., increase in subsystem entanglement) with increasing $L_{A}$.
Fluctuation in pSFF due to choice of subsystem :